OpenStudy (anonymous):

why doesnt the limit of 2x-12/|x+6| exist as you approach -6

OpenStudy (anonymous):

because if x=-6 and u input that in the absolute value ur gonna get 0. when u muiltiply the rest by 0 u get 0

OpenStudy (anonymous):

i guess i shouldnt have asked it like that because i understand that consept... my prof gave a hint and it says to evaluate the limit from the left and right

OpenStudy (anonymous):

Split the absolute value sign into two parts.

OpenStudy (anonymous):

\[ \frac{2x-12}{x+6}\quad \frac{2x-12}{-(x+6)} \]

OpenStudy (anonymous):

Suppose \[ \lim \frac{2x-12}{x+6} = L \]Then \[ \lim \frac{2x-12}{-(x+6)} = -L \]

OpenStudy (anonymous):

The only way for a limit to exist is if \(L = -L\).

OpenStudy (anonymous):

Or, basically it must be the case that \(L=0\).

OpenStudy (anonymous):

We can solve the limit using l'Hospital's rule: \[ \lim \frac{2x-12}{x+6} = \lim \frac{(2x-12)'}{(x+6)'}= \lim \frac{2}{1} = 2 \]

OpenStudy (anonymous):

Since \(L=2\neq 0\) the limit doesn't exist.